# Diffusion-Based Impedance Learning for Contact-Rich Manipulation Tasks

### v2

@@ Abstract
Learning methods excel at motion generation in the information domain but are not primarily designed for physical interaction in the energy domain. Impedance Control shapes physical interaction but requires task-aware tuning by selecting feasible impedance parameters. We present Diffusion-Based Impedance Learning, a framework that combines both domains. A Transformer-based Diffusion Model with cross-attention to external wrenches reconstructs a simulated Zero-Force Trajectory (sZFT). This captures both translational and rotational task-space behavior. For rotations, we introduce a novel SLERP-based quaternion noise scheduler that ensures geometric consistency. The reconstructed sZFT is then passed to an energy-based estimator that updates stiffness and damping parameters. A directional rule is applied that reduces impedance along non-task axes while preserving rigidity along task directions. Training data were collected for a parkour scenario and robotic-assisted therapy tasks using teleoperation with Apple Vision Pro. With only tens of thousands of samples, the model achieved sub-millimeter positional accuracy and sub-degree rotational accuracy. Its compact model size enabled real-time torque control and autonomous stiffness adaptation on a KUKA LBR iiwa robot. The controller achieved smooth parkour traversal within force and velocity limits and 30/30 success rates for cylindrical, square, and star peg insertions without any peg-specific demonstrations in the training data set. All code for the Transformer-based Diffusion Model, the robot controller, and the Apple Vision Pro telemanipulation framework is publicly available. These results mark an important step towards Physical AI, fusing model-based control for physical interaction with learning-based methods for trajectory generation.
@@ I Introduction
Robotic behavior emerges at the interface of two fundamentally different domains. Motion planning belongs to the information domain, where learning-based methods have recently shown remarkable progress. In contrast, physical interaction is governed by the energy domain, and model-based Impedance Control has been widely adopted to guarantee stability [1] and safety [2]. Robots that manipulate in unstructured environments must bridge these domains: motions must be inferred as information, yet executed through stable, energy-consistent exchange with the environment [3]. This requirement is critical in assembly, rehabilitation, and other contact-rich tasks where visual feedback is limited and success depends on regulating interaction rather than simply following a path.
@@
While Impedance Control provides the framework for shaping robot–environment interaction [4], its performance critically depends on selecting appropriate stiffness and damping parameters (which may need to be adapted online) [5]. Too much stiffness can lead to jamming, too little can prevent task execution.
Recent advances in contact-rich manipulation can be divided into two main strands. Model-based approaches incorporate explicit contact dynamics like friction cones, complementarity constraints, and mixed-integer formulations [6, 7, 8, 9, 10, 11, 12, 13, 14]. These methods produce behavior consistent with physical contact but face challenges: high computational cost, parameter sensitivity, and limited robustness in unstructured settings. Learning-based approaches, in contrast, operate in the information domain. Reinforcement learning and sampling-based optimization [15, 16, 17, 18], along with more recent generative models such as diffusion policies [19], flow-matching policies [20, 21], and Transformer-based policy models [22, 23], excel at motion generation across tasks and embodiments [24]. However, execution in these frameworks often relies on velocity-based or fixed-gain PD controllers. While fixed-gain PD controllers can avoid hardware damage, apparent compliance is only a byproduct of low control gains. In unstructured environments, where vision may be unreliable or occluded, explicit impedance regulation is essential to control interaction.
The Adaptive Compliance Policy of [25] is a first step toward combining generative models with compliance adaptation, but it is formulated within an admittance-control framework. Admittance control is known to struggle with transitions into and out of contact [4, 26], making it unsuitable for many assembly tasks such as peg-in-hole insertion. To our knowledge, no approach has successfully combined learning-based trajectory generation (information domain) with impedance adaptation (energy domain).
@@
This paper addresses that gap. We present Diffusion-Based Impedance Learning, a framework that unifies generative modeling with energy-consistent Impedance Control. A conditional Transformer-based Diffusion Model reconstructs simulated Zero-Force Trajectories [^1] The Zero-Force Trajectory (ZFT), introduced by Hogan [5], refers to the commanded equilibrium motion in the unconstrained case: the unique end-effector pose at which the interaction wrench vanishes. from contact-perturbed displacement and external wrench signals. The sZFT serves as a reconstructed equilibrium used to modulate task-space impedance. A directional adaptation scheme adapts stiffness primarily along non-task-relevant directions and preserves rigidity where needed for execution. In this way, our approach combines the strengths of learning-based methods (motion generation in the information domain) with those of model-based methods (impedance regulation in the energy domain).
@@
Fixed impedance $\times$ (stopped at first obstacle) 30/30 4/30 0/30
@@
We validate the framework on a KUKA LBR iiwa in two contact-rich scenarios: parkour-style obstacle traversal and multi-geometry peg-in-hole insertion. Training data were collected through telemanipulation, using Apple Vision Pro (AVP) equipped with a markerless pose tracking framework [27] integrated into the robot controller. Both tasks highlight the shortcomings of fixed impedance and simple adaptation schemes, which either jam at obstacles or fail to achieve insertion. In contrast, Diffusion-Based Impedance Learning achieves consistent success. The outcomes of Table I preview the central result of this work: bridging the information and energy domains is key to robust manipulation in unstructured environments.
All code for the Transformer-based Diffusion Model, the robot controller, and the Apple Vision Pro telemanipulation framework is available on our GitHub repository [^2] https://github.com/StrokeAIRobotics/DiffusionBasedImpedanceLearning. Demonstration videos of all experiments can be found on the project website [^3] https://strokeairobotics.github.io/DiffusionBasedImpedanceLearning.
II Methods
We outline the process in both domains: Transformer-based diffusion in the information domain and energy-based Impedance Control in the energy domain. These components are then fused through diffusion-based reconstruction of simulated Zero-Force Trajectories (sZFT).
@@ II-A Diffusion Models
Diffusion Models learn to sample from a data distribution by approximating the reverse-time dynamics of a Gaussian diffusion process. Let $\{z_{i}\}_{i=1}^{N}$ be independent and identically distributed samples from the data distribution $p_{\text{data}}$; we posit a parametric family of distributions $\{p_{\theta}\}_{\theta\in\Theta}$ and learn $\theta$ so that $p_{\theta}\approx p_{\text{data}}$. A forward Markov chain maps a clean sample $z_{0}$ to (approximately) $\mathcal{N}(0,\mathbf{I})$ over $T$ steps. The learned reverse-time chain iteratively denoises $z_{t}$ to produce a sample that approximates $p_{\text{data}}$ (Figure 1).
@@ II-A3 Reverse Process and Learning Objective
The true reverse $q(z_{t-1}\!\mid\!z_{t})$ is intractable, as it marginalizes over $z_{0}$. However, since the forward chain is linear–Gaussian, the posterior $q(z_{t-1}\!\mid\!z_{t},z_{0})$ is itself Gaussian with mean $\tilde{\mu}_{t}(z_{t},z_{0})$ and variance $\tilde{\beta}_{t}\mathbf{I}$ [31]. We model $p_{\theta}(z_{t-1}\!\mid\!z_{t})=\mathcal{N}(\mu_{\theta}(z_{t},t),\Sigma_{\theta}(z_{t},t))$ and fit $\theta$ by minimizing the variational bound $\sum_{t}D_{\mathrm{KL}}\!\big(q(z_{t-1}\!\mid\!z_{t},z_{0})\,\|\,p_{\theta}(z_{t-1}\!\mid\!z_{t})\big)$. With the $\varepsilon$-parameterization,
@@ II-A4 Training and Inference Algorithms
Training follows a noise-prediction objective: at each iteration a clean sample $z_{0}\!\sim\!p_{\text{data}}$, a timestep $t\!\in\!\{1,\dots,T\}$ (uniform), and Gaussian noise $\varepsilon\!\sim\!\mathcal{N}(0,\mathbf{I})$ are sampled. The closed-form forward diffusion produces $z_{t}$ and the network $\varepsilon_{\theta}$ is updated to minimize mean-squared error between $\varepsilon$ and its prediction $\varepsilon_{\theta}(z_{t},t)$ [30, 32]. At inference, generation starts from $z_{T}\!\sim\!\mathcal{N}(0,\mathbf{I})$ and applies the learned reverse transitions $p_{\theta}(z_{t-1}\!\mid\!z_{t})$ for $t{=}T,\dots,1$, yielding $z_{0}$ as a sample from the model [30]. For robotic trajectory generation, this iterative denoising maps pure noise into trajectories that are consistent with task goals (Figure 1).
@@ II-A5 Conditional Diffusion
In our approach, conditional diffusion is central. By conditioning on external wrenches, the model reconstructs trajectories that are consistent with physical interaction. This ties the denoising process in the information domain to the regulation of interaction in the energy domain.
More details about Diffusion Models can be found in [35, 30, 36, 32].
@@ II-A6 Transformer-based Diffusion Model
At each diffusion step, pose tokens are formed from translations and unit-quaternions. These tokens are embedded and enriched with learned positional and timestep embeddings. Interaction signals (forces and moments) are projected to context tokens and integrated through a multi-head cross-attention layer. Here, trajectory tokens serve as queries and context tokens as keys and values. The combined sequence is then processed by stacked self-attention and feed-forward layers, and a lightweight Multi-layer Perceptron head predicts the 7D noise (3D translations and 4D unit-quaterions) for each token. This architecture captures temporal structure while conditioning on interaction signals, a requirement that is central for contact-rich applications [37, 38, 39, 40, 41, 42].
More details about Transformer-based models can be found in [37, 41, 42, 36].
@@ II-B Energy-based Impedance Control
$\mathbf{e}_{r}$ $:=\log\!\big(\mathbf{q}_{0}\mathbf{q}^{-1}\big)\in\mathbb{R}^{3}\simeq\mathfrak{so}(3).$ (4c)
where the observed pose is $(\bm{p},\mathbf{q})\in\mathbb{R}^{3}\times\mathbb{S}^{3}$ and the equilibrium pose is $(\bm{p}_{0},\mathbf{q}_{0})\in\mathbb{R}^{3}\times\mathbb{S}^{3}$. Any unit quaternion admits the unit axis–angle representation
$\mathbf{q}=\big(\cos(\tfrac{\theta}{2}),\;\mathbf{u}\sin(\tfrac{\theta}{2})\big),\quad\|\mathbf{u}\|=1,\;\theta\in[0,\pi].$ (5)
In our formulation, the logarithm map is applied to the error quaternion $\mathbf{q}_{0}\mathbf{q}^{-1}$ to return the axis–angle error vector $\bm{e}_{r}=\bm{u}_{0}\ \theta_{0}$.
Let $\bm{K}_{t},\bm{K}_{r}\in\mathbb{R}^{3\times 3}$ be diagonal, positive semidefinite stiffness matrices and $\bm{K}=\mathrm{diag}(\bm{K}_{t},\bm{K}_{r})$. Define the potential energy
@@
The spatial wrench $\mathbf{F}=\begin{pmatrix}\mathbf{f}\\ \mathbf{m}\end{pmatrix}\in\mathfrak{se}^{\ast}(3)$ can obtained from the gradient
@@
$\mathbf{F}\;=\;\frac{\partial\mathcal{U}}{\partial\Delta\mathbf{x}}.$ (7)
The damping matrices $\bm{B}_{t},\bm{B}_{r}\in\mathbb{R}^{3\times 3}$ are chosen proportional to stiffness,
$\bm{B}_{t}=\lambda_{t}\bm{K}_{t},\qquad\bm{B}_{r}=\lambda_{r}\bm{K}_{r},$ (8)
with $\lambda_{t},\lambda_{r}>0$ selected as described in Appendix A-C. This yields Rayleigh-like dissipation that scales with stiffness and task-space inertia.
In addition to this virtual impedance, the environment contributes external wrenches $\bm{F}_{\text{ext}}\in\mathfrak{se}^{\ast}(3)$. These wrenches arise during both translational motion $\dot{\bm{p}}\in\mathbb{R}^{3}$ and rotational motion $\bm{\omega}\in\mathfrak{so}(3)$. Finally, the Jacobian $\bm{J}(\bm{q}):\mathbb{R}^{n}\rightarrow\mathfrak{se}(3)$, a function of the joint configuration $\bm{q}\in\mathbb{R}^{n}$, provides the mapping between the $n$-dimensional joint space and the 6-dimensional task space. An overview of the controller implemented in this paper can be found in Appendix A.
The Impedance Control framework can be illustrated by the Norton equivalent network, as represented in Figure 2. Here, two sources interact through the impedance element:
  • The motion source, which represents the equilibrium pose $\{\bm{p}_{0},\mathbf{q}_{0}\}$ at each point of the ZFT,
  • The impedance source, which represents the baseline stiffness $\bm{K}$. In our case, damping $\bm{B}$ is a function of stiffness and inertia (Appendix A-C).
The environment applies external wrenches $\bm{F}_{\text{ext}}$, which are balanced by the virtual spring–damper system. This ensures that the interaction wrenches are consistent both with the energy stored in the virtual spring and with the motion of the robot. This formulation highlights that Impedance Control operates not only in the information domain (motion planning through ZFT) but also in the energy domain (physical interaction through impedance), a key property that we exploit later for energy-based stiffness estimation (Section III-D).

### v3

@@ Abstract
Learning-based methods excel at robot motion generation but remain limited in contact-rich physical interaction. Impedance control provides stable and safe contact behavior but requires task-specific tuning of stiffness and damping parameters. We present Diffusion-Based Impedance Learning, a framework that bridges these paradigms by combining generative modeling with energy-consistent impedance control. A Transformer-based Diffusion Model, conditioned via cross-attention on measured external wrenches, reconstructs simulated Zero-Force Trajectories (sZFTs) that represent contact-consistent equilibrium behavior. A SLERP-based quaternion noise scheduler preserves geometric consistency for rotations on the unit sphere. The reconstructed sZFT is used by an energy-based estimator to adapt impedance online through directional stiffness and damping modulation. Trained on parkour and robot-assisted therapy demonstrations collected via Apple Vision Pro teleoperation, the model achieves sub-millimeter positional and sub-degree rotational accuracy using only tens of thousands of samples. Deployed in real-time torque control on a KUKA LBR iiwa, the approach enables smooth obstacle traversal and generalizes to unseen tasks, achieving 100% success in multi-geometry peg-in-hole insertion.
@@ I Introduction
Robotic behavior arises from the interaction of two fundamentally different domains. On the one hand, motion planning and learning algorithms are based on data processing and operate on symbolic or numerical representations of virtual desired behavior, such as trajectories, poses, or task objectives. These representations are processed unidirectionally and are constrained only by temporal causality (“no output before input”) and boundedness. They are not subject to physical laws. In this paper, we refer to this as the information domain [23].
Physical interaction with the environment, on the other hand, is governed by bidirectional behavior. It is constrained by physical laws like energy conservation, entropy production, gravity, etc. In this paper, we refer to this as the energy domain [23]. Ignoring these physical constraints will yield robot control problems with stability (passivity) [34, 36] and safety [32].
@@
Robots operating in contact-rich and unstructured environments must bridge these two domains: desired motions must be inferred and represented in the information domain, yet executed through control algorithms that take into account energy exchange with the environment. This requirement is critical in assembly, robot-assisted rehabilitation, and other contact-rich tasks where visual feedback may be limited and task success depends on regulating physical interaction rather than simply following a path.
The relation between motion planning (information domain) and physical interaction (energy domain) is well described by mechanical impedance [19, 38, 63]. Impedance Control has been widely adopted to guarantee stability [19] and safety [32] for robots that go in and out of contact. The separation of both domains can be formalized by equivalent networks, a generalization of the equivalent-circuit concept [13], to represent nonlinear dynamics and arbitrarily complex physical networks as a combination of primitives [21, 50, 51].
@@
While Impedance Control provides a powerful framework for shaping robot–environment interaction [21], its effectiveness critically depends on the choice of stiffness and damping parameters [17]. Too much stiffness can lead to jamming, while too little can prevent task execution. Selecting these parameters is typically an iterative and time-consuming process, as current physics-based simulation tools are not yet sufficiently accurate or reliable to support fully offline tuning [11].
Recent advances in contact-rich manipulation can be divided into three main strands: 1) model-based approaches, 2) statistical-impedance heuristics, and 3) learning-based approaches.
Model-based approaches incorporate explicit contact dynamics including friction cones, complementarity constraints, and mixed-integer formulations [56, 45, 54, 46, 16, 4, 5, 66, 65, 40, 41, 58, 3, 47]. These methods produce behavior consistent with physical contact but face challenges: high computational cost, parameter sensitivity, and limited robustness in unstructured settings. Most approaches require task-specific tuning of parameters. The selected control gains typically do not generalize across environments and tasks. Consequently, these approaches are not ideal for unstructured environments with unknown contact dynamics. These limitations motivate the development of learning-based frameworks capable of inferring and adapting stiffness online from observed interaction data.
Between explicit physics-based models and fully learning-based policies, a class of statistical-impedance heuristics have emerged within the learning-from-demonstration literature. These approaches typically model demonstrated motions using probabilistic representations such as Gaussian Mixture Models, where the mean trajectory defines the nominal motion and the covariance is interpreted as a measure of task relevance or confidence [1, 52]. Impedance parameters are then modulated heuristically based on this statistical variability, with higher stiffness assigned to directions of low variance. These approaches have shown to be effective for structured tasks with consistent geometry. However, these methods are still limited in contact-rich settings where changing environmental constraints are not reflected in the statistical variability of the demonstrations.
Learning-based approaches, in contrast, primarily operate in the information domain. Reinforcement learning and sampling-based optimization methods [69, 29, 64, 73, 75, 67] can achieve context-aware behavior, but typically lack an explicit representation of impedance, which prescribes a bidirectional relationship between motion and generalized force [18]. Instead, these methods rely on reward functions that penalize either force or position deviations. As a result, they tend to converge toward admittance-like behaviors that suppress interaction forces or motion deviations, rather than explicitly regulating impedance.
More recent generative approaches, including diffusion policies [9], flow-matching policies [7, 76], and Transformer-based policy models [60, 30], have demonstrated impressive performance in motion generation across tasks and embodiments [68]. However, these methods primarily focus on trajectory generation and are typically executed through PD controllers with fixed proportional and derivative gains to maintain stability during contact. Compliance in such systems is therefore not explicitly controlled but emerges indirectly through low control gains. While this strategy can prevent hardware damage, it provides limited control over physical interaction behavior in contact-rich or safety-critical scenarios where deliberate impedance modulation is essential.
Recent work has begun to combine generative models with compliant control for contact-rich manipulation. The Adaptive Compliance Policy (ACP)[25] represents a way toward integrating generative models with compliance adaptation, where target poses are inferred from visual observations and tracked during execution. ACP has demonstrated smooth behavior in visuomotor tasks such as object manipulation and surface interaction. However, ACP is formulated within an admittance-control framework, which is known to be sensitive to transitions into and out of contact[21, 39]. This sensitivity limits its applicability in contact-rich assembly tasks such as peg-in-hole insertion, where stable regulation of contact forces is essential.
@@
This paper addresses this gap by formulating impedance adaptation as an equilibrium reconstruction problem rather than a trajectory learning problem. We present Diffusion-Based Impedance Learning, a framework that integrates generative modeling with energy-consistent impedance control. While the robot moves on a fixed nominal Zero-Force Trajectory (ZFT) [^1] The Zero-Force Trajectory (ZFT), introduced by Hogan [17], refers to the commanded equilibrium motion in the unconstrained case: the unique end-effector pose at which the interaction wrench vanishes., a conditional Transformer-based Diffusion Model reconstructs a simulated Zero-Force Trajectory (sZFT) from contact-perturbed motion and measured interaction wrenches. The reconstructed sZFT represents an inferred equilibrium that is consistent with the observed physical interaction and serves as the reference for impedance modulation. A directional adaptation scheme selectively reduces stiffness along non-task-relevant directions while preserving rigidity along directions essential for task execution.
In this way, our approach bridges data-driven learning and energy-consistent control: Diffusion Models explain physical interaction by reconstructing contact-consistent equilibria, while impedance control enforces stability and physical consistency during execution. Our method operates purely on robot kinematics and measured interaction wrenches, without relying on visual observations.
We validated the framework on a KUKA LBR iiwa in two contact-rich scenarios: parkour-style obstacle traversal and multi-geometry peg-in-hole insertion. Training data were collected through teleoperation using Apple Vision Pro (AVP) equipped with a markerless pose-tracking framework [55] integrated into the robot controller. Both tasks expose the limitations of fixed impedance and simple adaptation schemes, which require careful, task-specific parameter tuning and still tend to jam at obstacles or fail to achieve insertion. In contrast, Diffusion-Based Impedance Learning achieves consistent success without manual retuning.
This result is particularly noteworthy in the peg-in-hole experiments, as the Diffusion Model was not trained on this task. Despite the absence of peg-in-hole demonstrations during training, the system achieved a 100% success rate across all peg geometries. This generalization was obtained using only  1.6 hours of teleoperated training data (72,000 samples). The results summarized in Table I underscore the central contribution of this work: robust contact-rich manipulation in unstructured environments requires explicitly bridging learning in the information domain with control in the energy domain.
@@
Fixed impedance $\times$ (stopped at first obstacle) 100% 13% 0%
@@
The code for the Transformer-based Diffusion Model, the robot controller, and the Apple Vision Pro teleoperation framework is available in our GitHub repository [^2] https://GitHub.com/diffusion-imp-anon/DiffusionBasedImpedanceLearning. Demonstration videos of all experiments can be found on the project website [^3] https://diffusion-imp-anon.GitHub.io/DiffusionBasedImpedanceLearning/.
II Preliminaries
First, we outline the process in both domains: Transformer-based diffusion in the information domain and energy-based Impedance Control in the energy domain.
@@ II-A Diffusion Models
Diffusion Models learn to sample from a data distribution by approximating the reverse-time dynamics of a Gaussian diffusion process [28]. Let $\{z_{i}\}_{i=1}^{N}$ be independent and identically distributed samples from the data distribution $p_{\text{data}}$; the goal is to learn a parametric family of distributions $\{p_{\theta}\}_{\theta\in\Theta}$ such that $p_{\theta}\approx p_{\text{data}}$. A forward Markov chain maps a clean sample $z_{0}$ to (approximately) $\mathcal{N}(0,\mathbf{I})$ over $T$ steps. The learned reverse-time chain iteratively denoises $z_{t}$ to produce a sample that approximates $p_{\text{data}}$ (Figure 1).
@@ II-A3 Reverse Process and Learning Objective
The true reverse $q(z_{t-1}\!\mid\!z_{t})$ is intractable, as it marginalizes over $z_{0}$. However, since the forward chain is linear–Gaussian, the posterior $q(z_{t-1}\!\mid\!z_{t},z_{0})$ is itself Gaussian with mean $\tilde{\mu}_{t}(z_{t},z_{0})$ and variance $\tilde{\beta}_{t}\mathbf{I}$ [61]. The model represents $p_{\theta}(z_{t-1}\!\mid\!z_{t})=\mathcal{N}(\mu_{\theta}(z_{t},t),\Sigma_{\theta}(z_{t},t))$ and fit $\theta$ by minimizing the variational bound $\sum_{t}D_{\mathrm{KL}}\!\big(q(z_{t-1}\!\mid\!z_{t},z_{0})\,\|\,p_{\theta}(z_{t-1}\!\mid\!z_{t})\big)$. With the $\varepsilon$-parameterization,
@@ II-A4 Training and Inference Algorithms
Training follows a noise-prediction objective: at each iteration, a clean sample $z_{0}\!\sim\!p_{\text{data}}$, a timestep $t\!\in\!\{1,\dots,T\}$ (uniform), and Gaussian noise $\varepsilon\!\sim\!\mathcal{N}(0,\mathbf{I})$ are sampled. The closed-form forward diffusion produces $z_{t}$, then the network $\varepsilon_{\theta}$ is updated to minimize mean-squared error between $\varepsilon$ and its prediction $\varepsilon_{\theta}(z_{t},t)$ [15, 24]. At inference, generation starts from $z_{T}\!\sim\!\mathcal{N}(0,\mathbf{I})$ and applies the learned reverse transitions $p_{\theta}(z_{t-1}\!\mid\!z_{t})$ for $t{=}T,\dots,1$, yielding $z_{0}$ as a sample from the model [15]. For robotic trajectory generation, this iterative denoising maps pure noise into trajectories that are consistent with task goals (Figure 1).
@@ II-A5 Conditional Diffusion
In our approach, conditional diffusion is central. By conditioning on external wrenches, the model reconstructs sZFTs that are consistent with physical interaction. This ties the denoising process in the information domain to the regulation of interaction in the energy domain. More details about Diffusion Models can be found in [62, 15, 74, 24].
@@ II-A6 Transformer-based Diffusion Model
At each diffusion step, pose tokens are formed from translations and unit-quaternions. These tokens are embedded and enriched with learned positional and timestep embeddings. Interaction signals (forces and moments) are projected to context tokens and integrated through a multi-head cross-attention layer. Here, trajectory tokens serve as queries while context tokens serve as keys and values. The combined sequence is then processed by stacked self-attention and feed-forward layers, and a lightweight Multi-layer Perceptron (MLP) head predicts the 7D noise (3D translations and 4D unit-quaternions) for each token. This architecture captures temporal structure while conditioning on interaction signals, a requirement that is central for contact-rich applications [70, 42, 12, 27, 71, 43]. More details about Transformer-based models can be found in [70, 71, 43, 74].
@@ II-B Energy-based Impedance Control
$\mathbf{e}_{r}$ $:=\log\!\big(\bm{Q}_{0}\bm{Q}^{-1}\big)\in\mathbb{R}^{3}\simeq\mathfrak{so}(3).$ (4c)
where the observed pose is $(\bm{p},\bm{Q})\in\mathbb{R}^{3}\times\mathbb{S}^{3}$ and the equilibrium pose is $(\bm{p}_{0},\bm{Q}_{0})\in\mathbb{R}^{3}\times\mathbb{S}^{3}$. Any unit quaternion admits the unit axis–angle representation
$\bm{Q}=\big(\cos(\tfrac{\theta}{2}),\;\mathbf{u}\sin(\tfrac{\theta}{2})\big),\quad\|\mathbf{u}\|=1,\;\theta\in[0,\pi].$ (5)
In this formulation, the logarithm map is applied to the error quaternion $\bm{Q}_{0}\bm{Q}^{-1}$ to return the axis–angle error vector $\bm{e}_{r}=\bm{u}_{0}\ \theta_{0}$.
Let $\bm{K}_{t},\bm{K}_{r}\in\mathbb{R}^{3\times 3}$ be diagonal, positive semidefinite translational and rotational stiffness matrices and $\bm{K}=\mathrm{diag}(\bm{K}_{t},\bm{K}_{r})$. Define the potential energy
@@
The elastic wrench $\mathbf{F}_{\text{elastic}}\in\mathfrak{se}^{\ast}(3)$ can then be derived from the gradient
@@
$\mathbf{F}_{\text{elastic}}\;=\;\frac{\partial\mathcal{U}}{\partial\Delta\mathbf{x}}=\begin{pmatrix}\bm{f}_{\text{elastic}}\\[2.0pt] \bm{m}_{\text{elastic}}\end{pmatrix}$, (7)
with $\bm{f}_{\text{elastic}}\in\mathbb{R}^{3}$ and $\bm{m}_{\text{elastic}}\in\mathbb{R}^{3}$ being the elastic force and moment, respectively.
II-B1 Damping design
The damping design was formulated to ensure that energy dissipation scales consistently with both the commanded stiffness $\bm{K}_{\text{\{t,r}\}}$ and the task-space inertia for translations and rotations $\bm{\Lambda}_{\text{\{t,r}\}}\in\mathbb{R}^{3\times 3}$. To account for numerical issues near singular configurations, $\bm{\Lambda}_{\text{\{t,r}\}}$ was computed using a damped least-squares method [34].
$\bm{D}_{\text{\{t,r}\}}=d\,\bm{I}_{3}.$ (8)
The symmetric positive semidefinite square roots $\sqrt{\bm{\Lambda}_{\text{\{t,r}\}}}$ and $\sqrt{\bm{K}_{\text{\{t,r}\}}}$ are obtained by eigendecomposition for $\bm{\Lambda}_{\text{\{t,r}\}}$ and elementwise for $\bm{K}_{\text{\{t,r}\}}$. As presented in [2], an intermediate damping matrix $\bm{B}_{\text{\{t,r}\}}$ is constructed as
$\bm{B}_{\text{\{t,r\}}}$ $=\sqrt{\bm{\Lambda}_{\text{\{t,r\}}}}\,\bm{D}_{\text{\{t,r\}}}\,\sqrt{\bm{K}_{\text{\{t,r\}}}}$ (9)
$\quad+\sqrt{\bm{K}_{\text{\{t,r\}}}}\,\bm{D}_{\text{\{t,r\}}}\,\sqrt{\bm{\Lambda}_{\text{\{t,r\}}}}.$
From $\bm{B}_{\text{\{t,r}\}}$, a scalar damping coefficient is computed to normalize the dissipation relative to the overall stiffness:
$\lambda_{\text{\{t,r}\}}\;=\;\frac{2\,\mathrm{tr}(\bm{B}_{\text{\{t,r}\}})}{\mathrm{tr}(\bm{K}_{\text{\{t,r}\}})}\;=\;\frac{2\,\mathrm{tr}(\bm{B}_{\text{\{t,r}\}})}{k_{x}+k_{y}+k_{z}}.$ (10)
Finally, the damping matrices can be expressed as
$\bm{B}_{t}=\lambda_{t}\,\bm{K}_{t},$ (11)
for translation and
$\bm{B}_{r}=\lambda_{r}\,\bm{K}_{r},$ (12)
for rotation.
II-B2 Control Force
The force $\bm{f}_{\text{ctrl}}\in\mathbb{R}^{3}$ is composed of $\bm{f}_{\text{elastic}}$ and damping through matrix $\bm{B}_{t}\in\mathbb{R}^{3\times 3}$ that acts on the end-effector translational velocity $\dot{\bm{p}}\in\mathbb{R}^{3}$. It can be calculated as
$\bm{f}_{\text{ctrl}}=\bm{f}_{\text{elastic}}-\bm{B}_{t}\ \dot{\bm{p}}.$ (13)
II-B3 Control Moment
The moment $\bm{m}_{\text{ctrl}}\in\mathbb{R}^{3}$ is composed of $\bm{m}_{\text{elastic}}$ and damping through matrix $\bm{B}_{r}\in\mathbb{R}^{3\times 3}$ that acts on the angular velocity of the end-effector $\bm{\omega}\in\mathfrak{so}(3)$. It can be calculated as
$\bm{m}_{\text{ctrl}}=\bm{m}_{\text{elastic}}-\bm{B}_{r}\ \bm{\omega}.$ (14)